OpenStudy (anonymous):

Find the centroid of the region bounded by the given curves. y = x^3, x + y = 10, y = 0

OpenStudy (anonymous):

I have found the area to be 36.

OpenStudy (anonymous):

Yet, my x,y values for the centroid are incorrect.

OpenStudy (accessdenied):

How did you calculate your x and y values? This can possibly help us determine where you went wrong.

OpenStudy (anonymous):

For x: \[1/A \int\limits_{0}^{2} x(10-x-x^3)dx\] For y: \[1/A \int\limits_{0}^{2} 1/2 (10-x-x^3)^2dx\]

OpenStudy (accessdenied):

What happened to the region from 2 to 10? This was only from 0 to 2. :) I think we will have to add up two integrals, one for the section [0,2] and one for the section [2,10]. y = x^3 makes up the upper boundary of the region over [0,2], and y= -x + 10 makes the upper bound of the region over [2, 10]. Adding them together ensures that you are getting contribution from each section in the final answer.

OpenStudy (anonymous):

Oh, gosh. So for x: \[1/36 [\int\limits_{0}^{2}(x^3-(-x+10))dx + \int\limits_{2}^{10}(-x+10-x^3)dx ]\] ?

OpenStudy (accessdenied):

The piece of the region on the interval [0,2] is bounded from above by x^3 and below by y=0. The piece of the region on the interval [2, 10] is bounded from above by y = -x + 10 and below by y=0. So f(x) = x^3 and g(x) = 0 for [0,2] and f(x) = -x+10 and g(x) = 0 for the other.

OpenStudy (anonymous):

For x: \[1/36[\int\limits_{0}^{2}(x^3)dx+\int\limits_{2}^{10}(-x+10)dx]\]

OpenStudy (accessdenied):

Almost there. I think you just missed a small detail. That setup looks like the area for the region, but I believe we need a little extra multiplied on the integrand for these coordinates.

OpenStudy (anonymous):

Multiply by x?

OpenStudy (accessdenied):

Yes. These would be called moments of the region. They are already set up that way if you are using the formulas. http://tutorial.math.lamar.edu/Classes/CalcII/CenterOfMass.aspx For the x-coordinate, we are multiplying both integrands by x before we integrate. The y-coordinate is somewhat different but we apply the same strategy suggested by the formula in both integrals to find the y-coordinate.

ganeshie8 (ganeshie8):

if you're familiar wid double integrals, \(\large \overline{x} = \frac{1}{A}\iint x dA\)

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

wat value u getting for x ?

OpenStudy (anonymous):

Can you stick around so im sure my setup for y is right?

ganeshie8 (ganeshie8):

ahh never saw those formulas before, they look cool :)

OpenStudy (accessdenied):

Hmm.. I haven't seen it approached like that. :o Nice! And yes, I can stick around. :)

OpenStudy (anonymous):

Ok, umm for y: \[1/72[\int\limits_{0}^{2}(x^3)^2dx+\int\limits_{2}^{10}(-x+10)^2dx]\]

OpenStudy (accessdenied):

I think that looks good as well.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (accessdenied):

glad to help! plotting all the graph information together, this is what turned out: http://www.wolframalpha.com/input/?i=Plot+y+%3D+x%5E3%2C+y+%3D+-x%2B10%2C++y+%3D+0%2C+x%3D+4.3%2C+y+%3D+2.6%3B+from+y%3D0+to+y%3D10 i think the point looks fairly accurate to where i'd place the center of mass.

OpenStudy (anonymous):

My online hw says the answer is incorrect

OpenStudy (accessdenied):

what was the input?

OpenStudy (anonymous):

584/135, 332/135

OpenStudy (accessdenied):

hm.. i think there was an error evaluating the y-value's integral? wolfram computes the integral to be 496/189. http://www.wolframalpha.com/input/?i=%28integral+from+0+to+2+of+%28x%5E3%29%5E2+dx+%2B+integral+from+2+to+10+of+%2810-x%29%5E2+dx%29%2F72

OpenStudy (accessdenied):

if you have the work written down, you could look it over step-by-step to make sure there wasn't any fishy business in the middle.

OpenStudy (anonymous):

Thanks, I have the work. I made an error.

OpenStudy (anonymous):

I have the correct answer now, thank you so much again!

OpenStudy (accessdenied):

Yup, you're welcome! :)